find 2f(x) - f(x-1) for f(x)=3x^2 + 5
Answer by edjones(8007)

2f(x)=2(3x^2+5)=6x^2+10
.
f(x-1)=3(x-1)^2+5
=3(x^2-2x+1)+5
=3x^2-6x+3+5
=3x^2-6x+8
.
6x^2+10-(3x^2-6x+8)
=6x^2+10-3x^2+6x-8
=3x^2+6x+2
.
Ed